OpenStudy (anonymous):

find slope intercept for x-8y=8 with points (2,4)

OpenStudy (anonymous):

y=1/8x-1

OpenStudy (anonymous):

so the slope is 1/8

OpenStudy (anonymous):

can you show me how you got this plz

OpenStudy (anonymous):

Slope intercept form: y=mx+b ......... which m being the slope and b being the y intercept x-8y=8 -8y= -x+8 8y= x - 8 y=1/8x-1

OpenStudy (anonymous):

so y is 1/8 and x is -1

OpenStudy (anonymous):

@artishaalston what are you asking to find?

OpenStudy (anonymous):

slope intercept of the perpendicular to the line x-8y=8 with 2,4

OpenStudy (anonymous):

slope perpindicular to the equation x-8y=8?

OpenStudy (anonymous):

yes using 2,4

OpenStudy (anonymous):

2-8(4)y=8

OpenStudy (anonymous):

ok what does this mean

OpenStudy (anonymous):

wait nvm mind hold up

OpenStudy (anonymous):

4=8(2)+b 4=16+b -16 -16 4=b

OpenStudy (anonymous):

srry B=12

OpenStudy (anonymous):

-12

OpenStudy (anonymous):

B is your new Y

OpenStudy (precal):

just like stated earlier